OpenStudy (muscrat123):

WILL MEDAL AND FAN (06.02)A bag contains 30 marbles, 3 of which are blue, 20 are red, 5 are green, and 2 are purple. Emerson takes a marble out of the bag, records the color, and returns it to the bag. How many purple marbles should he expect after 300 trials? 20 30 50 200

OpenStudy (muscrat123):

@Hero @Preetha

OpenStudy (muscrat123):

anybody?

OpenStudy (muscrat123):

@nerlineg

OpenStudy (anonymous):

On any draw, what is the probability of choosing a purple marble?

OpenStudy (muscrat123):

idk, but i think it is 20

OpenStudy (anonymous):

How many purple marbles are there?

OpenStudy (kropot72):

The probability of drawing a purple marble is 2/30 on each trial. The expected number of purple marbles after 300 trials is therefore \[\large \frac{2}{30}\times\frac{300}{1}=?\]

OpenStudy (muscrat123):

i think so because all of the marbles added up = 30, and 30 x 10 = 300

OpenStudy (muscrat123):

2 x 10 = 20

OpenStudy (kropot72):

Correct.

OpenStudy (muscrat123):

and 600/30

OpenStudy (muscrat123):

so it is 20?

OpenStudy (kropot72):

Yes.

OpenStudy (muscrat123):

thank u!

OpenStudy (kropot72):

you're welcome :)